# Metaplectic time-frequency representations

### v2

@@ Abstract.
Time-frequency representations stemmed in 1932 with the introduction of the Wigner distribution. For most of the 20th century, research in this area primarily focused on defining joint probability distributions for position and momentum in quantum mechanics. Applications to electrical engineering were soon established with the seminal works of Gabor and the researchers at Bell Labs. In 2012, Bai, Li and Cheng used for the first time metaplectic operators, defined in the middle of 20th century by Van Hove, to generalize the Wigner distribution and unify effectively the most used time-frequency representations under a common framework. This work serves as a comprehensive up-to-date survey on time-frequency representations defined by means of metaplectic operators, with particular emphasis on the recent contributions by Cordero and Rodino, who exploited metaplectic operators to their limits to generalize the Wigner distributions. Their idea provides a fruitful framework where properties of time-frequency representations can be explained naturally by the structure of the symplectic group.
@@ 2.1. The 20th century
[31], that was named after Rihacek due to his contribution in 1968 [40]. Parallel to these works, Gabor proposed time-frequency shifts
@@
for some $\sigma\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$. Besides the Wigner distribution, the ambiguity function, and the Rihacek distribution, other elements of the Cohen’s class are discussed in [5, 33, 36, 48].
@@
where $0\leq\tau\leq 1$. Observe that for $\tau=0$, $W_{\tau}=W_{0}$ is the Rihacek distribution, whereas the classical Wigner distribution corresponds to the choice $\tau=1/2$. The cross-$\tau$-Wigner distributions were considered in [3].
@@ 2.2. The metaplectic framework
In 2015, Z. Zhang and Luo generalized the new Wigner–Ville transform further [47], improving its ability of detecting linear frequency modulated signals under low Signal-to-Noise Ratio. Their construction can be formalized in dimension $d\geq 1$ as follows: fixed $\hat{S},\hat{S}_{1},\hat{S}_{2}\in\operatorname*{Mp}(d,\mathbb{R})$, and $\mathfrak{T}_{M_{1/2}}f(x,t)=f(x+t/2,x-t/2)$,
@@
Parallel to the works by Z. Zhang et al, Bayer, Cordero, Gröchenig and Trapasso defined the (cross-)matrix Wigner distributions in 2019 [2, 19]. They observed that a peculiar detail shared by the distributions in (8), (5), (9) and (10) is that they can be written as
@@ 2.3. Metaplectic Wigner distributions
    then $W_{\hat{S},\hat{S}_{1},\hat{S}_{2}}(f,g)=\widehat{\mathcal{A}_{\hat{S},\hat{S}_{1},\hat{S}_{2}}}(f\otimes\bar{g})$, where $W_{\hat{S},\hat{S}_{1},\hat{S}_{2}}(f,g)$ is the cross-version of (14), and
@@ 3.2. Shift-invertiblity
In [18] the authors conjectured condition $E_{\mathcal{A}}\in\operatorname*{GL}(d,\mathbb{R})$ to be crucial in the identification of those metaplectic Wigner distributions so that (31) holds.
@@
    for some constant $c_{\mathcal{A}}\in\mathbb{C}$, $|c_{\mathcal{A}}|=1$, where $\widehat{\delta_{\mathcal{A}}}=\mathcal{F}\widehat{\overline{G_{\mathcal{A}}}}$.
@@
Theorem 3.6 tells that shift-invertibility is sufficient, along with other minor conditions, to characterize modulation spaces. On the contrary, the Rihacek distribution is not shift-invertible with
@@
and right-hand side is not proportional to $\lVert f\rVert_{M^{p,q}}$, unless $p=q=2$. This fact is more general. In [26] it is proved that $M^{p}$ quasi-norms cannot be measured by means of non shift-invertible distributions.
@@ 4. Inversion formula and Gabor frames
(46) $f\overset{L^{2}}{=}\frac{1}{\langle\gamma,g\rangle}\int_{{\mathbb{R}^{2d}}}W_{g}f(z)\pi_{\mathcal{A}}(z)\gamma dz.$
In general, metaplectic atoms map $\mathcal{S}(\mathbb{R}^{d})$ to $\mathcal{S}^{\prime}(\mathbb{R}^{d})$. However, metaplectic atoms of shift-invertible metaplectic Wigner distributions define surjective quasi-isometries of $L^{2}(\mathbb{R}^{d})$. In this case, we can consider a system $\mathcal{G}_{\mathcal{A}}(g,\Lambda)=\{\pi_{\mathcal{A}}(\lambda)g\}_{\lambda\in\Lambda}$, where $g\in L^{2}(\mathbb{R}^{d})\setminus\{0\}$ and $\Lambda\subset{\mathbb{R}^{2d}}$ is discrete. We say that $\mathcal{G}_{\mathcal{A}}$ is a metaplectic Gabor frame (for $L^{2}(\mathbb{R}^{d})$) if it is a frame for $L^{2}(\mathbb{R}^{d})$. In view of (40), it is straightforward to relate metaplectic Gabor frames to classical Gabor frames, see [13, Theorem 6.4]. In the case of shift-invertible distributions, the deformation operator $\widehat{\delta_{\mathcal{A}}}$ in (40) plays a central role in the frame theory associated to $W_{\mathcal{A}}$. The frame operator associated to $\mathcal{G}_{\mathcal{A}}(g,\Lambda)$ is $S_{\mathcal{A}}f=\sum_{\lambda\in\Lambda}W_{\mathcal{A}}(f,g)(\lambda)\pi_{\mathcal{A}}(\lambda)g$, $f\in L^{2}(\mathbb{R}^{d})$, and the canonical dual window of $g$ is $\gamma_{\mathcal{A}}=\widehat{\delta_{\mathcal{A}}}^{-1}S_{\mathcal{A}}^{-1}\widehat{\delta_{\mathcal{A}}}g$. We have that if $m\asymp m\circ E_{\mathcal{A}}$, then for every $0<p,q\leq\infty$,

### v3

@@ Abstract.
Time-frequency representations stemmed in 1932 with the introduction of the Wigner distribution. For most of the 20th century, research in this area primarily focused on defining joint quasi-probability distributions for position and momentum in quantum mechanics. Applications to electrical engineering were soon established with the seminal works of Gabor and the researchers at Bell Labs. In 2012, Bai, Li and Cheng used for the first time metaplectic operators, defined in the middle of 20th century by Van Hove, to generalize the Wigner distribution and unify effectively the most used time-frequency representations under a common framework. This work serves as a comprehensive up-to-date survey on time-frequency representations defined by means of metaplectic operators, with particular emphasis on the recent contributions by Cordero and Rodino, who exploited metaplectic operators to their limits to generalize the Wigner distributions. Their idea provides a fruitful framework where properties of time-frequency representations can be explained naturally by the structure of the symplectic group.
@@ 2.1. The 20th century
[31], that was named after Rihaczek due to his contribution in 1968 [40]. Parallel to these works, Gabor proposed time-frequency shifts
@@
for some $\sigma\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$. Besides the Wigner distribution, the ambiguity function, and the Rihaczek distribution, other elements of the Cohen’s class are discussed in [5, 33, 36, 48].
@@
where $0\leq\tau\leq 1$. Observe that for $\tau=0$, $W_{\tau}=W_{0}$ is the Rihaczek distribution, whereas the classical Wigner distribution corresponds to the choice $\tau=1/2$. The cross-$\tau$-Wigner distributions were considered in [3].
@@ 2.2. The metaplectic framework
In 2015, Zhang and Luo generalized the new Wigner–Ville transform further [47], improving its ability of detecting linear frequency modulated signals under low Signal-to-Noise Ratio. Their construction can be formalized in dimension $d\geq 1$ as follows: fixed $\hat{S},\hat{S}_{1},\hat{S}_{2}\in\operatorname*{Mp}(d,\mathbb{R})$, and $\mathfrak{T}_{M_{1/2}}f(x,t)=f(x+t/2,x-t/2)$,
@@
Parallel to the works by Zhang et al, Bayer, Cordero, Gröchenig and Trapasso defined the (cross-)matrix Wigner distributions in 2019 [2, 19]. They observed that a peculiar detail shared by the distributions in (8), (5), (9) and (10) is that they can be written as
@@ 2.3. Metaplectic Wigner distributions
    then $W_{\hat{S},\hat{S}_{1},\hat{S}_{2}}(f,g)=\widehat{\mathcal{A}_{\hat{S},\hat{S}_{1},\hat{S}_{2}}}(f\otimes\bar{g})$, where $W_{\hat{S},\hat{S}_{1},\hat{S}_{2}}(f,g)$ is the polarized version of (14), and
@@ 3.2. Shift-invertiblity
In [18] the authors conjectured condition $E_{\mathcal{A}}\in\operatorname*{GL}(2d,\mathbb{R})$ to be crucial in the identification of those metaplectic Wigner distributions so that (31) holds.
@@
    for some constant $c_{\mathcal{A}}\in\mathbb{C}$, $|c_{\mathcal{A}}|=1$, where $\widehat{\delta_{\mathcal{A}}}=\mathcal{F}\widehat{\overline{G_{\mathcal{A}}}}$, where ${\widehat{\overline{G_{\mathcal{A}}}}}f=\widehat{{G_{\mathcal{A}}}}\bar{f}$.
@@
Theorem 3.6 tells that shift-invertibility is sufficient, along with other minor conditions, to characterize modulation spaces. On the contrary, the Rihaczek distribution is not shift-invertible with
@@
and right-hand side is not proportional to $\lVert f\rVert_{M^{p,q}}$, unless $p=q=2$. This fact is more general. In [26] it is proved that $M^{p}$ cannot be characterized by means of non shift-invertible distributions.
@@ 4. Inversion formula and Gabor frames
(46) $f\overset{L^{2}}{=}\frac{1}{\langle\gamma,g\rangle}\int_{{\mathbb{R}^{2d}}}W_{\mathcal{A}}(f,g)(z)\pi_{\mathcal{A}}(z)\gamma dz.$
In general, metaplectic atoms map $\mathcal{S}(\mathbb{R}^{d})$ to $\mathcal{S}^{\prime}(\mathbb{R}^{d})$. However, metaplectic atoms of shift-invertible metaplectic Wigner distributions define surjective quasi-isometries of $L^{2}(\mathbb{R}^{d})$. In this case, we can consider a system $\mathcal{G}_{\mathcal{A}}(g,\Lambda)=\{\pi_{\mathcal{A}}(\lambda)g\}_{\lambda\in\Lambda}$, where $g\in L^{2}(\mathbb{R}^{d})\setminus\{0\}$ and $\Lambda\subset{\mathbb{R}^{2d}}$ is discrete. We say that $\mathcal{G}_{\mathcal{A}}(g,\Lambda)$ is a metaplectic Gabor frame (for $L^{2}(\mathbb{R}^{d})$) if it is a frame for $L^{2}(\mathbb{R}^{d})$. In view of (40), it is straightforward to relate metaplectic Gabor frames to classical Gabor frames, see [13, Theorem 6.4]. In the case of shift-invertible distributions, the deformation operator $\widehat{\delta_{\mathcal{A}}}$ in (40) plays a central role in the frame theory associated to $W_{\mathcal{A}}$. The frame operator associated to $\mathcal{G}_{\mathcal{A}}(g,\Lambda)$ is $S_{\mathcal{A}}f=\sum_{\lambda\in\Lambda}W_{\mathcal{A}}(f,g)(\lambda)\pi_{\mathcal{A}}(\lambda)g$, $f\in L^{2}(\mathbb{R}^{d})$, and the canonical dual window of $g$ is $\gamma_{\mathcal{A}}=\widehat{\delta_{\mathcal{A}}}^{-1}S_{\mathcal{A}}^{-1}\widehat{\delta_{\mathcal{A}}}g$. We have that if $m\asymp m\circ E_{\mathcal{A}}$, then for every $0<p,q\leq\infty$,